P(A) = 0.5
P(B) = 0.6
P(A and B) = ?

Found 2 solutions by ikleyn, math_tutor2020:
Answer by ikleyn(52780)

.

Impossible to answer.


As given, the problem is INCOMPLETE and the question CAN NOT be answered.



Answer by math_tutor2020(3816)


If events A and B are independent, then we can say:
P(A and B) = P(A)*P(B)
P(A and B) = 0.5*0.6
P(A and B) = 0.3

If A and B were dependent, then we cannot use that formula.
Instead we would have to say something like
P(A and B) = P(A)*P(B given A)
or
P(A and B) = P(B)*P(A given B)
Many math textbooks will replace the word "given" with a vertical bar.

Since the instructions do not make it clear if A & B are independent or not, there is unfortunately not enough information to answer this problem.